### v2

# Some Geometric Properties of the Yff Points of a Triangle International Journal of Computer Discovered Mathematics (IJCDM) ISSN 2367-7775 ©IJCDM Volume 11, 2026 pp. xx–$y$

Abstract. The Yff points of a triangle were introduced by Peter Yff in 1963. Since then, very few new facts have been discovered about these points. We present some geometrical properties of the Yff points of various shaped triangles which were discovered and proved by computer.
Keywords. bicentric pairs, computer-discovered mathematics,Yff points.
Mathematics Subject Classification (2020). 51M04, 51-08.
1. Introduction
The following result was found by Peter Yff in 1963 [8].
Theorem $1.$
Let $D_{1}$, $D_{2}$, $E_{1}$, $E_{2}$, $F_{1}$, $F_{2}$ be points on the sides of triangle ABCABC (as shown in Figure 1) such that $AF_{1}=BD_{1}=CE_{1}=AE_{2}=BF_{2}=CD_{2}=u$ where $u$ is the real root of the equation $x^{3}=(a-x)(b-x)(c-x)$. Then $AD_{1}$, $BE_{1}$, $CF_{1}$ meet in a point $Y_{1}$ and $AD_{2}$, $BE_{2}$, $CF_{2}$ meet in a point $Y_{2}$.
Figure 1. Yff Points
The points $Y_{1}$ and $Y_{2}$ have become known as the Yff Points of a triangle. If $\triangle ABC$ is named counterclockwise, then $Y_{1}$ is the 1st Yff point and $Y_{2}$ is the 2nd Yff point.
Very little is known about these points that was not found in Yff’s original paper in 1963. It is clear from the definition that the Yff points are isotomic conjugates. A summary of what is known about Yff points can be found in [8], [7], [9, §3.3.2], and [6].
The value $u$ is connected to $a$, $b$, and $c$ by the equation
(1) $u^{3}=(a-u)(b-u)(c-u).$
The value of $u$ can be expressed in terms of radicals as follows.
$u=\frac{\sqrt[3]{k_{1}}}{6\sqrt[3]{2}}-\frac{k_{4}}{3\ 2^{2/3}\sqrt[3]{k_{1}}}+\frac{1}{6}(a+b+c)$
where
$k_{1}$ $=k_{3}+\sqrt{k_{2}},$
$k_{2}$ $=k_{3}^{2}+4k_{4}^{3},$
$k_{3}$ $=2(a+b+c)^{3}-18(ab+bc+ca)(a+b+c)+108abc,$
and
$k_{4}$ $=6(ab+ac+bc)-(a+b+c)^{2}.$
This representation is useful when constructing the Yff points in Dynamic Geometry Enviroments such as Geometer’s Sketchpad or GeoGebra.
Yff also found the following interesting result, where $X_{n}$ denotes the $n$-th named triangle center in the Encyclopedia of Triangle Centers [5].
Theorem 1.2.
Let ABCABC be an arbitrary triangle (Figure 2). Then $Y_{1}Y_{2}\perp X_{1}X_{3}$.
Figure 2. red lines are perpendicular
Open Question 1.
Is there a purely synthetic geometric proof of Theorem 1.2 (without using coordinates)?
In this paper, we will give some new geometric properties of Yff points in various shape triangles.
We found these properties using Mathematica. We started with a random numerical triangle. As $n$ and $m$ ranged from 1 to 30, we looked at the triangle plus the Yff points plus $X_{n}$ and $X_{m}$ and checked (numerically) for various properties in the resulting figure, such as whether two lines were parallel, two segments had equal length, two angles had equal measure, etc. We did not find any results that were not equivalent to Theorem 1.2.
We did the same thing for various shaped triangles, such as right triangles, heptagonal triangles, and triangles whose sides are in arithmetic progression. The results are presented in the next sections. Since these results were found numerically, valid to 10 decimal places, the computations did not constitute a proof that the result was true. At this point, we then used Mathematica again, this time using exact symbolic computations to prove formally that each result was true.
2. Right Triangles
Theorem 2.1.
Let ABCABC be a right triangle (named counterclockwise) with right angle at $B$ (Figure 3). Then $AY_{2}$, $CY_{1}$, and $X_{8}X_{20}$ are concurrent.
Figure 3. dashed lines are concurrent
Proof.
Set up a barycentric coordinate system using $\triangle ABC$ as the reference triangle, so that he coordinates for AA, $B$, and $C$ are $(1:0:0)$, $(0:1:0)$, and $(0:0:1)$, respectively. The coordinates for $X_{8}$ and $X_{20}$ can be found in [5].
$X_{8}$ $=a-b-c::$
$X_{20}$ $=3a^{4}-2a^{2}b^{2}-2a^{2}c^{2}-b^{4}+2b^{2}c^{2}-c^{4}::$
where “f(a,b,c)f(a,b,c)::” denotes the coordinates $\bigl(f(a,b,c):f(b,c,a):f(c:a:b)\bigr)$.
The coordinates for $Y_{1}$ and $Y_{2}$ can be found in [8].
$Y_{1}$ $=\left(\frac{c-u}{b-u}\right)^{1/3}::$
$Y_{2}$ $=\left(\frac{b-u}{c-u}\right)^{1/3}::$
where $u$ is given in Theorem $1.$ We will call these the symmetric coordinates for $Y_{1}$ and $Y_{2}$.
Yff also gives simpler coordinates for $Y_{1}$ that are not as symmetric: $\left(\frac{u}{b-u}:\frac{a-u}{u}:1\right)$. Eliminating fractions and finding similar coordinates for $Y_{2}$ give
(2) $Y_{1}$ $=\left(u^{2}:(a-u)(b-u):u(b-u)\right)$
$Y_{2}$ $=\left((a-u)(b-u):u^{2}:u(a-u)\right)$
We will call these the simple coordinates for $Y_{1}$ and $Y_{2}$.
If we represent the line $px+qy+rz=0$ by the triple $(p:q:r)$, then, we get the nice symmetric properties that the line through two points is given in Mathematica by the vector cross product Cross[point1,point2] and the point of intersection of two lines is given by Cross[line1,line2]. Having found the barycentric coordinates for all the points in Figure 3, we can now calculate the equations of the three dashed lines. To avoid cube roots, we use the coordinates for $Y_{1}$ and $Y_{2}$ as given in display (2) .
From [9, §4.3], we have that three lines $(p_{i}:q_{i}:r_{i})$ are concurrent if and only if
$\left|\begin{array}[]{lll}p_{1}&q_{1}&r_{1}\\ p_{2}&q_{2}&r_{2}\\ p_{3}&q_{3}&r_{3}\end{array}\right|=0.$
In our case, the condition for the three dashed lines to be concurrent is
(3) $-2a^{7}bu+2a^{7}u^{2}+2a^{6}b^{2}u-4a^{6}bcu+4a^{6}bu^{2}+4a^{6}cu^{2}-6a^{6}u^{3}+4a^{5}b^{3}u$
$-4a^{5}b^{2}u^{2}+6a^{5}bcu^{2}-4a^{5}bu^{3}-6a^{5}cu^{3}+4a^{5}u^{4}-4a^{4}b^{4}u-4a^{4}b^{3}u^{2}$
$+6a^{4}b^{2}u^{3}+4a^{4}bc^{3}u-4a^{4}bc^{2}u^{2}-2a^{4}bcu^{3}+4a^{4}bu^{4}-4a^{4}c^{3}u^{2}+4a^{4}c^{2}u^{3}$
$-2a^{3}b^{5}u+6a^{3}b^{4}u^{2}+2a^{3}bc^{4}u-4a^{3}bc^{3}u^{2}+4a^{3}bc^{2}u^{3}-2a^{3}c^{4}u^{2}+4a^{3}c^{3}u^{3}$
$-8a^{3}c^{2}u^{4}+2a^{2}b^{6}u+4a^{2}b^{5}cu-4a^{2}b^{4}cu^{2}-2a^{2}b^{4}u^{3}-4a^{2}b^{3}c^{3}u-2a^{2}b^{2}c^{4}u$
$+4a^{2}b^{2}c^{3}u^{2}+2a^{2}c^{4}u^{3}-4ab^{6}u^{2}-6ab^{5}cu^{2}+4ab^{5}u^{3}+4ab^{4}c^{2}u^{2}+6ab^{4}cu^{3}$
$-4ab^{4}u^{4}+8ab^{3}c^{3}u^{2}-4ab^{3}c^{2}u^{3}-8ab^{2}c^{3}u^{3}-2abc^{5}u^{2}+2ac^{5}u^{3}+4ac^{4}u^{4}$
$+2b^{6}u^{3}+2b^{5}cu^{3}-4b^{5}u^{4}-4b^{4}c^{2}u^{3}-4b^{3}c^{3}u^{3}+8b^{3}c^{2}u^{4}+2b^{2}c^{4}u^{3}+2bc^{5}u^{3}$
$-4bc^{4}u^{4}=0$
Note that the radical form for $u$ is too complex, so we have left $u$ as an unevaluated variable, noting that $u$ satisfies Equation (1).
We can eliminate $u$ from equations (3) and (1). This can be accomplished using elimination theory and Gröbner bases. The process is straightforward, but tedious and is best left for computers. In Mathematica, the Eliminate function is used for this purpose. Eliminating $u$ from equations (3) and (1) gives
$a^{11}b^{4}c^{2}-a^{10}b^{6}c+3a^{10}b^{5}c^{2}-2a^{10}b^{4}c^{3}+a^{9}b^{7}c+3a^{9}b^{5}c^{3}-a^{9}b^{4}c^{4}+2a^{8}b^{8}c$
$5a^{8}b^{7}c^{2}+a^{8}b^{6}c^{3}-2a^{7}b^{9}c+a^{7}b^{8}c^{2}+a^{7}b^{4}c^{6}-a^{6}b^{10}c+3a^{6}b^{9}c^{2}-a^{6}b^{6}c^{5}$
$3a^{6}b^{5}c^{6}+2a^{6}b^{4}c^{7}+a^{5}b^{11}c-3a^{5}b^{9}c^{3}-a^{5}b^{8}c^{4}+5a^{5}b^{7}c^{5}-3a^{5}b^{5}c^{7}+a^{5}b^{4}c^{8}$
$a^{4}b^{11}c^{2}+a^{4}b^{10}c^{3}+2a^{4}b^{9}c^{4}-2a^{4}b^{8}c^{5}-a^{4}b^{7}c^{6}+a^{4}b^{6}c^{7}=0$
Polynomials of degree 17 in three variables are not easy to factor, but using the Factor function in Mathematica, we find that this condition is equivalent to
$a^{4}b^{4}c(a-c)(a-b+c)(a+b+c)(a^{2}-b^{2}+c^{2})(a^{2}c+ab^{2}-3abc+ac^{2}-b^{3}+b^{2}c)=0.$
The factor $a^{2}-b^{2}+c^{2}$ shows that this condition will be true for a right triangle with vertex at $B$. ∎
This proof is not satisfying. The computation involved in this proof is enormous. It seems highly unlikely that Theorem 2.1 is true merely because the determinant condition for collinearity happens to evaluate to 0. The points $Y_{1}$ and $Y_{2}$ have nice geometric definitions as shown in Figure 1. The points $X_{8}$ and $X_{20}$ also have nice geometric definitions. (Point $X_{8}$ is the Nagel Point of a triangle and Point $X_{20}$, also known as the de Longchamps Point of a triangle, is the reflection of the orthocenter about the circumcenter.)
Open Question 2.
Is there a purely synthetic geometric proof of Theorem 2.1?
Theorem $2.$
Let ABCABC be a right triangle (named counterclockwise) with right angle at $B$ and $\angle ACB=30^{\circ}$ (Figure 4). Then $AY_{1}$, $BY_{2}$, and $X_{8}X_{21}$ are concurrent.
Figure 4. dashed lines are concurrent
Proof.
Using the same procedure as in the proof of Theorem 2.1, we find that the condition for concurrence is $(a-b)(a^{4}+a^{3}b-2a^{2}c(b+c)+ab(b^{2}-2bc-3c^{2})+(b^{2}-c^{2})^{2})(a+b-c-2u)(a-u)(b-u)u=0.$ Eliminating $u$ from this equation and Equation (1) gives $a^{4}(a-b)b^{4}c(a+b+c)(a^{2}b+ab^{2}-3abc+ac^{2}+bc^{2}-c^{3})(a^{3}+b^{3}-a^{2}c-abc-b^{2}c-ac^{2}-bc^{2}+c^{3})=0.$ Setting $b=\sqrt{a^{2}+c^{2}}$ and $a=c\sqrt{3}$ shows that this condition is equivalent to $0=0$ which is always true. ∎
The following result comes from [9, §4.5].
Lemma 2.3.
The infinite point on the line represented by $(p:q:r)$ is (qr:rp:pq)(q-r:r-p:p-q).
Theorem 2.4.
Let ABCABC be a right triangle with right angle at $B$ (Figure 5). Then $BX_{8}\perp Y_{1}Y_{2}$.
Figure 5. blue lines are perpendicular
Proof.
[…]
Since $BX_{8}$ and $X_{1}X_{3}$ have the same point at infinity, they must be parallel. But $X_{1}X_{3}\perp Y_{1}Y_{2}$ by Theorem 1.2. Therefore, $BX_{8}\perp Y_{1}Y_{2}$. ∎
3. $60^{\circ}$ Triangles
Theorem 3.1.
Let ABCABC be a triangle (named counterclockwise) with $\angle C=60^{\circ}$ (Figure 6). Then $AY_{1}$, $BY_{2}$, and $X_{3}X_{8}$ are concurrent.
Figure 6. dashed lines are concurrent
Proof.
As before, we find that the condition for concurrence is
$(a-b)(a+b-c)(a^{2}-ab+b^{2}-c^{2})(a+b-c-2u)u=0.$
From the Law of Cosines, we see that in this triangle, $c^{2}=a^{2}+b^{2}-ab$, so this condition holds. ∎
4. Heptagonal Triangles
A heptagonal triangle is a triangle whose angles are $\pi/7$, $2\pi/7$, and $4\pi/7$.
Lemma 4.1.
Let ABCABC be a heptagonal triangle with $a>b>c$. Then
$a:b:c=\sin\frac{4\pi}{7}:\sin\frac{2\pi}{7}:\sin\frac{\pi}{7}.$
Proof.
This follows immediately from the Law of Sines. ∎
Lemma 4.2 (Reflection About BC).
Let $P$ have barycentric coordinates $(p:q:r)$ with respect to $\triangle ABC$. Then $Q$, the reflection of $P$ about side BCBC, has coordinates
$\left(a^{2}p:(c^{2}-b^{2})p-a^{2}(p+q):(b^{2}-c^{2})p-a^{2}(p+r)\right).$
Proof.
The formula for the distance between two points in barycentric coordinates can be found in [9, §7.1]. Let $Q=(p^{\prime}:q^{\prime}:r^{\prime})$. Solving the two equations $PB=QB$ and $PC=QC$ for $p^{\prime}$, $q^{\prime}$, and $r^{\prime}$ and excluding the solution where $Q=P$ gives the required coordinates for $Q$. ∎
Another proof, not using the distance formula, is instructive.
Proof.
It is immediate that infinite point of the line through the points $(p_{1}:q_{1}:r_{1})$ and $(p_{2}:q_{2}:r_{2})$ with the same sum $p_{1}+q_{1}+r_{1}=p_{2}+q_{2}+r_{2}$ is given by $(p_{1}-p_{2}:q_{1}-q_{2}:r_{1}-r_{2})$. Thus, if $(p,q,r)$ is a given point, we can consider the point $X=(0:qS_{B}+(p+q)S_{C}:(p+r)S_{B}+rS_{C})$ with sum $(p+q+r)(S_{B}+S_{C})=(p+q+r)a^{2}$. Then, by subtraction
$\left({0:q{S_{B}}+(p+q){S_{C}}:r{S_{C}}+(p+r){S_{B}}}\right)-{a^{2}}\left({p:q:r}\right)$
$=$ $\left({-{a^{2}}p:q{S_{B}}+(p+q){S_{C}}-{a^{2}}q:r{S_{C}}+(p+r){S_{B}}-{a^{2}}r}\right)$
$=$ $\left({-{a^{2}}p:p{S_{C}}:p{S_{B}}}\right),$
that is, the infinite point of PXPX is the infinite point of the AA-altitude and then PXPX is perpendicular to BCBC.
Again, by subtraction,
$2\left({0:q{S_{B}}+(p+q){S_{C}}:r{S_{C}}+(p+r){S_{B}}}\right)-{a^{2}}\left({p:q:r}\right)$
$=$ $\left({-{a^{2}}p:q\cdot 2{S_{B}}+(p+q)\cdot 2{S_{C}}-{a^{2}}q:r\cdot 2{S_{C}}+(p+r)\cdot 2{S_{B}}-{a^{2}}r}\right)$
$=$ $\left({-{a^{2}}p:2{S_{C}}p+{a^{2}}q:2{S_{B}}p+{a^{2}}r}\right),$
therefore the reflection of $P$ in $X$ is the point
$\left({-{a^{2}}p:({b^{2}}-{c^{2}})p+(p+q){a^{2}}:({c^{2}}-{b^{2}})p+(p+r){a^{2}}}\right).\qed$
Theorem 4.3.
Let ABCABC be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$ (Figure 7). Let $Y_{1}^{\prime}$ be the reflection of $Y_{1}$ about side BCBC. Then $X_{1}Y_{2}\parallel Y_{1}^{\prime}C$.
Figure 7. red lines are parallel
Proof.
From Lemma 4.2, we find that the coordinates of $Y_{1}^{\prime}$ are
$\left(a^{2}u^{2}:a^{2}(b-2u)u+(-b^{2}+c^{2})u^{2}+a^{3}(-b+u):u(-a^{2}b+(b^{2}-c^{2})u)\right).$
Using Lemma 2.3, we find that the infinite point on line $CY_{1}^{\prime}$ is
$i_{1}=\Bigl(a^{2}u^{2}:a^{2}(b-2u)u+(-b^{2}+c^{2})u^{2}+a^{3}(-b+u)$
$a^{3}(b-u)+(b^{2}-c^{2})u^{2}+a^{2}u(-b+u)\Bigr).$
Similarly, the infinite point on the line $X_{1}Y_{2}$ is
$i_{2}=\Bigl(a(b+c)(b-u)-a^{2}u-(b+c)(b-u)u:u(b^{2}+cu)+a(-b^{2}+u^{2})$
$a^{2}u+a(b-u)(-c+u)+u(b(c-u)-2cu)\Bigr)$
[…]
$a=\sin\frac{4\pi}{7},\,b=\sin\frac{2\pi}{7},\,\mathrm{and}\,\,c=\sin\frac{\pi}{7}.$
Thus, the lines are parallel. ∎
Theorem $4.$
Let ABCABC be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$ (Figure 8). Then $\angle CX_{1}Y_{2}+\angle X_{1}CY_{1}=\frac{\pi}{7}$.
Figure 8.
Proof.
Let $Y_{1}^{\prime}$ be the reflection of $Y_{1}$ about BCBC. Draw in line $CX_{3}$ as shown in Figure 9.
Figure 9.
Since $X_{3}$ is the center of the circle through AA, $B$, and $C$, $\angle CX_{3}B=6\pi/7$. This implies $\angle BCX_{3}=\pi/14$. Since $X_{1}$ is the incenter, $CX_{1}$ bisects $\angle ACB$, so $\angle X_{1}CB=\pi/14$. Therefore $\angle BCX_{3}=\angle X_{1}CB$.
Since $Y_{1}^{\prime}$ is the reflection of $Y_{1}$ about side BCBC, $\angle Y_{1}CB=\angle BCY_{1}^{\prime}$, i.e. the dotted angles are equal (the “black angles”). Subtracting, we find that $\angle X_{1}CY_{1}=\angle Y_{1}^{\prime}CX_{3}$, i.e. the green angles are equal.
From Theorem 4.3, we know that $Y_{2}X_{1}\parallel CY_{1}^{\prime}$. Therefore orange = green + 2 black. Hence orange + green = 2 green + 2 black = 2(green + black) = $2\angle X_{1}CB=\angle ACB=\pi/7$. ∎
[…]
Theorem 4.5.
Let ABCABC be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$. Let the angles be named as shown in Figure 10. If $\angle Y_{2}X_{1}A=\theta$, then
$\angle 1$ $=\frac{2\pi}{7}$
$\angle 2$ $=\frac{9\pi}{14}-\theta$
$\angle 3$ $=\frac{\pi}{14}$
$\angle 4$ $=\theta-\frac{\pi}{2}$
$\angle 5$ $=\frac{11\pi}{14}$
$\angle 6$ $=\frac{4\pi}{7}-\theta$
$\angle 7$ $=\frac{\pi}{7}$
$\angle 8$ $=\frac{\pi}{7}$
$\angle 9$ $=\frac{2\pi}{7}$
Figure 10.
Proof.
Since $\triangle ABC$ is a heptagonal triangle, $\angle CBA=\pi/7$, $\angle ACB=2\pi/7$, and $\angle BAC=4\pi/7$. Since $X_{1}$ is the incenter of $\triangle ABC$, $AX_{1}$, $BX_{1}$, and $CX_{1}$ are angle bisectors. Thus, $\angle 1=\angle 9=2\pi/7$, $\angle 3=\angle 4+\angle 6=\pi/14$, and $\angle 7=\angle 8=\pi/7$. From $\triangle CX_{1}A$, we find $\angle 2=9\pi/14-\theta$. From $\triangle BX_{1}C$, we find $\angle 5=11\pi/14$. From Theorem 4.4, $\angle 2+\angle 4=\pi/7$. Therefore $\angle 4=\pi/7-(9\pi/14-\theta)=\theta-\pi/2$. Finally, $\angle 6=\angle X_{1}CB-\angle 4=\pi/14-(\theta-\pi/2)=4\pi/7-\theta$. ∎
Theorem 4.6.
Let ABCABC be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$ (Figure 11). Then $\angle CX_{1}Y_{2}=\angle Y_{1}CX_{3}$.
Figure 11. green angles are equal
Proof.
Since $X_{3}$ is the center of the circle through AA, $B$, and $C$, $\angle CX_{3}B=6\pi/7$. This implies $\angle BCX_{3}=\pi/14$. Then, using Theorem 4.5, we have $\angle Y_{1}CX_{3}=\angle BCX_{3}+\angle 6=\pi/14+4\pi/7-\theta=9\pi/14-\theta=\angle 2$. ∎
We also found the following three results, involving other centers. The proofs, using Mathematica, involved a lot of computation, so we have omitted the details.
Theorem 4.7.
Let ABCABC be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$. Let $P$ be the foot of the perpendicular from $Y_{1}$ to ACAC (Figure 12). Then $\angle CBX_{5}=\angle Y_{2}Y_{1}P$.
Figure 12. green angles are equal
Theorem 4.8.
Let ABCABC be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$ (Figure 13). Then $\angle X_{21}X_{1}Y_{2}=\angle CY_{1}Y_{2}$.
Figure 13. green angles are equal
Theorem 4.9.
Let ABCABC be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$ (Figure 14). Then $\angle X_{1}Y_{2}Y_{1}=\angle X_{28}CY_{1}$.
Figure 14. green angles are equal
Open Question 3.
Is there a purely synthetic geometric proof of any of the three preceding theorems, or a proof that does not involve massive amounts of computation?
5. Harmonic Triangles
Theorem 5.1.
Let ABCABC be a triangle such that $Y_{1}Y_{2}$ passes through a vertex. Then $Y_{1}Y_{2}$ is a median of the triangle.
Proof.
Without loss of generality, assume that $Y_{1}Y_{2}$ passes through vertex B and $\triangle ABC$ is named counterclockwise (Figure 15).
Figure 15. $B$, $Y_{1}$, and $Y_{2}$ are collinear
Let $BY_{1}$ meet ACAC at $M$. Then, from Figure 1 (left), we have $CM=u$. Line $BY_{2}$ is the same line, so also meets ACAC at $M$. From Figure 1 (right), we have $AM=u$. Thus, $AM=CM$ and BMBM is a median of the triangle. ∎
Theorem 5.2.
Let ABCABC be a triangle such that $BY_{1}$ is a median. Then $Y_{1}$ divides the median into two segments in the ratio 2b:2ab2b:2a-b.
Proof.
Let $BY_{1}$ meet ACAC at $E$ and let $AY_{1}$ meet BCBC at $D$.
Then, from Figure 1 we see that $CE=BD=u$. Since $BY_{1}$ is a median, $AE=CE$, so $AE=u$. Since $BD=u$, we have CD=auCD=a-u. Line ADAD is a transversal of $\triangle EBC$, so by the Theorem of Menelaus, we have x(au)u=yu(2u)x(a-u)u=yu(2u), so x/y=2u/(au)x/y=2u/(a-u). But $u=b/2$. Thus $x/y=b/(a-\frac{b}{2})$ which implies x/y=2b/(2ab)x/y=2b/(2a-b). ∎
Corollary 5.3.
If $Y_{1}$ is the midpoint of a median, then the triangle is similar to a 3–4–6 triangle.
Theorem 5.4.
Let ABCABC be a triangle such that $BY_{2}$ is a median. Then $Y_{2}$ divides the median into two segments in the ratio 2(2ab):b2(2a-b):b.
Proof.
The proof is similar to the proof of Theorem 5.2, so the details are omitted. ∎
The harmonic mean of $x$ and $y$ is $\frac{2}{1/x+1/y}$. A triangle is called a harmonic triangle if one side is the harmonic mean of the other two sides. We say that the triangle is harmonic.
Theorem $5.$
A triangle is harmonic if and only if $Y_{1}Y_{2}$ passes through a vertex.
Proof.
If $Y_{1}Y_{2}$ passes through a vertex, assume without loss of generality, that the vertex is $B$ and $\triangle ABC$ is named counterclockwise (Figure 15). From Theorem 5.1, $BY_{1}$ is a median and so $u=b/2$. Letting $u=b/2$ in Equation (1) gives b(ab2ac+bc)=0b(ab-2ac+bc)=0 which implies $b=2ac/(a+c)$ and the triangle is harmonic.
Conversely, if the triangle is harmonic, assume without loss of generality that $2/b=1/a+1/c$. The barycentric coordinates for $B$ are $(0:1:0)$. Using the simple coordinates for $Y_{1}$ and $Y_{2}$, we find that that $Y_{1}Y_{2}$ passes through $B$ if and only if
$\left|\begin{array}[]{ccc}u^{2}&(a-u)(b-u)&(b-u)u\\ (a-u)(b-u)&u_{2}&(a-u)u\\ 0&1&0\end{array}\right|=0.$
Expanding this gives b(b2u)(au)u=0b(b-2u)(a-u)u=0. From [8], it is known that $0<u<a$, so therefore, $Y_{1}Y_{2}$ passes through vertex $B$ in a harmonic triangle, if and only if $b=2u$.
If the triangle is harmonic, then $b=2u$, so bu=ub-u=u, and from Figure 1, $BY_{1}$ is a median. Similarly, $BY_{2}$ is a median. Therefore $Y_{1}Y_{2}$ passes through $B$. ∎
Theorem 5.6.
A triangle is harmonic if and only if $Y_{1}Y_{2}$ passes through $X_{2}$.
Proof.
Assume that $Y_{1}Y_{2}$ passes through $X_{2}$. The barycentric coordinates for $X_{2}$ are $(1:1:1)$. Using the simple coordinates for $Y_{1}$ and $Y_{2}$, the determinant condition for $Y_{1}$, $Y_{2}$ and $X_{2}$ to be collinear is
$\left|\begin{array}[]{ccc}u^{2}&(a-u)(b-u)&(b-u)u\\ (a-u)(b-u)&u_{2}&(a-u)u\\ 1&1&1\end{array}\right|=0.$
Expanding this out gives (a2u)(b2u)(buab+au)=0(a-2u)(b-2u)(bu-ab+au)=0. Eliminating variable $u$ between this equation and Equation (1) gives
$a^{3}b^{3}(ab+ac-2bc)(2ab-ac-bc)(ab-2ac+bc)=0$
which shows that $\triangle ABC$ is harmonic.
Conversely, if the triangle is harmonic, then by Theorem 5.5, $Y_{1}Y_{2}$ passes through a vertex. From Theorem 5.1, we can conclude that $Y_{1}Y_{2}$ is a median. Thus, $Y_{1}Y_{2}$ passes through $X_{2}$. ∎
Theorem 5.7.
A triangle is harmonic if and only if $Y_{1}Y_{2}$ bisects a side of the triangle.
Proof.
Without loss of generality. assume that $Y_{1}Y_{2}$ passes through $E$, the midpoint of ACAC. The barycentric coordinates for $E$ are $(1:0:1)$. Using the simple coordinates for $Y_{1}$ and $Y_{2}$, the determinant condition for $Y_{1}$, $Y_{2}$ and $E$ to be collinear is
$\left|\begin{array}[]{ccc}u^{2}&(a-u)(b-u)&(b-u)u\\ (a-u)(b-u)&u_{2}&(a-u)u\\ 1&0&1\end{array}\right|=0.$
Expanding this out gives $(b-2u)(a^{2}(b-u)+2au(u-b)+bu^{2})=0$. Eliminating variable $u$ between this equation and Equation (1) gives
(4) $a^{3}b^{4}(ab-2ac+bc)(a^{2}b+2a^{2}c-2abc+2ac^{2}+bc^{2})=0.$
If $a^{2}b+2a^{2}c-2abc+2ac^{2}+bc^{2}=0$, then $b=-2ac(a+c)/(a-c)^{2}$. This equation is impossible since the left side is positive and the right side is negative. Thus ab2ac+bc=0ab-2ac+bc=0 which implies that the triangle is harmonic.
Conversely, if the triangle is harmonic, then by Theorem 5.5, $Y_{1}Y_{2}$ passes through a vertex. From Theorem 5.1, we can conclude that $Y_{1}Y_{2}$ is a median. Thus, $Y_{1}Y_{2}$ passes through $E$. ∎
6. AP Triangles
An AP Triangle is a triangle whose sides are in arithmetic progression.
Theorem 6.1.
Let $\triangle ABC$ (named counterclockwise), be an AP Triangle, with $2a=b+c$. Then $AX_{1}\parallel Y_{1}Y_{2}$.
Figure 16. red lines are parallel
Proof.
According to [6], the point at infinity on $Y_{1}Y_{2}$ is $X_{513}$. From [5], we find that the barycentric coordinates of $X_{513}$ are
$I_{1}=\left(a(b-c):b(c-a):c(a-b)\right).$
Since $A=(1:0:0)$ and $X_{1}=(a:b:c)$, the point at infinity on $AX_{1}$ is
$I_{2}=(-b-c:b:c)$
When b=2acb=2a-c, we find that $I_{1}=(-2a:2a-c:c)$ and $I_{2}=(-2a:2a-c:c)$. Thus, the two lines have the same point at infinity. Therefore $AX_{1}\parallel Y_{1}Y_{2}$. ∎
7. Double-Angle Triangles
A double-angle triangle is a triangle in which one angle is twice another angle. Note that all heptagonal triangles are double-angle triangles.
The following result is fairly well known.
Lemma 7.1 (Double Angle Condition).
Let ABCABC be a triangle in which $\angle A=2\angle B$. Then $a^{2}=b(b+c)$.
Proof.
Let $\angle CBA=\theta$, so that $\angle BAC=2\theta$. Extend CACA past AA to a point $D$ such that $AD=AB$, so that ABDABD is an isosceles triangle and $\angle ABD=\angle BDA$.
By the Exterior Angle Theorem, $\angle BAC=\angle ABD+\angle BDA$. Thus, $\angle ABD=\angle BDA=\theta$ and $\triangle BAC\sim\triangle DBC$. Hence $AC/BC=BC/DC$ or $b/a=a/(b+c)$.
Theorem 7.2.
Let ABCABC be a double-angle triangle (named counterclockwise) with $\angle A=2\angle B$ (Figure 17). Let $Y_{2}^{\prime}$ be the reflection of $Y_{2}$ about side BCBC. Then $BY_{2}$, $CX_{2}$, and $AY_{2}^{\prime}$ are concurrent.
Figure 17. dashed lines are concurrent
Proof.
From Lemma 4.2, we find that the coordinates of $Y_{2}^{\prime}$ are
$\Bigl(a^{2}(a-u)(b-u):(-b^{2}+c^{2})(a-u)(b-u)+a^{2}((b-2u)u+a(-b+u))$
$(-a^{2}b+(b^{2}-c^{2})(b-u))(a-u)\Bigr).$
We can therefore write down the equations for the three lines $AY_{2}^{\prime}$, $BY_{2}$, and $CX_{2}$. The determinant condition for these three lines to be concurrent simplifies to
$(a-u)(-ab(b^{2}-c^{2})(b-u)+a^{3}(b-u)^{2}+b(b^{2}-c^{2})(b-u)u-a^{2}u(b^{2}-2bu+2u^{2}))=0.$
Eliminating variable $u$ between this equation and Equation (1) gives
$ab(b-c)(a^{2}-b(b+c))=0.$
Thus, the three lines are concurrent if $a^{2}=b(b+c)$. The theorem now follows from Lemma 7.1. ∎
8. Isosceles Triangles
Theorem 8.1.
Let $Y$ be one of the Yff points of $\triangle ABC$. Let the areas of triangles BCYBCY, CAYCAY, and ABYABY be $K_{a}$, $K_{b}$, and $K_{c}$, respectively. Then $K_{b}K_{c}=K_{a}^{2}$ if and only if $\triangle ABC$ is isosceles with $AB=AC$.
Proof.
Start with $Y=Y_{1}$. The areas of these triangles are proportional to the barycentric coordinates of $Y_{1}$. Then from Equation (2), without loss of generality, we have $K_{a}=u^{2}$, $K_{b}=(a-u)(b-u)$, and $K_{c}=(b-u)u$. Thus, $K_{b}K_{c}=K_{a}^{2}$ is equivalent to
(5) $(a-u)u(-b+u)^{2}=u^{4}.$
Eliminating $u$ from Equations (5) and (1), this equation becomes $a^{4}b^{4}(b-c)c=0$. This condition is true if and only if $b=c$. The proof when $Y=Y_{2}$ is similar. ∎
9. Isosceles Right Triangles
The following two results were found by computer. We omit the complicated proofs.
Theorem 9.1.
Let ABCABC be an isosceles right triangle (named counterclockwise) with right angle at AA. Let $X_{1}Y_{2}$ meet BCBC at $D$ (Figure 18). Then $\angle DY_{2}C=\angle CBX_{8}$.
Figure 18. green angles are equal
Theorem 9.2.
Let ABCABC be an isosceles right triangle (named counterclockwise) with right angle at AA. Then $\angle Y_{2}X_{1}C=\angle Y_{2}CX_{9}$ (Figure 19).
Figure 19. green angles are equal
10. Special Triangles
Theorem 10.1.
In $\triangle ABC$ (named counterclockwise), $BY_{1}$ passes through $X_{8}$ if and only if
$a^{2}(a-b-c)=bc(b+c-3a).$
Figure 20 shows an example.
Figure 20. $B$, $Y_{1}$, and $X_{8}$ are collinear
Proof.
Let $BX_{8}$ meet ACAC at $E$. Then by a well-known property of the Nagel point, CE=saCE=s-a where $s$ is the semiperimter of the triangle [9, §3.2.2].
Then $Y_{1}$ lies on BEBE, if and only if sa=us-a=u. Eliminating $u$ between this condition and Equation (1) gives the condition to be $a^{2}(a-b-c)=bc(b+c-3a)$ as required. ∎
In the same manner, we get the following theorem.
Theorem 10.2.
In $\triangle ABC$ (named counterclockwise), $CY_{1}$ passes through $X_{7}$ if and only if
$a^{2}(a-b-c)=bc(b+c-3a).$
11. Cevian Triangles
Let $[XYZ]$ denote the area of $\triangle XYZ$.
Lemma 11.1 (The Common Angle Theorem).
In $\triangle ABC$, let $D$ be a point on BCBC and let $E$ be a point on ACAC (Figure 21). Then
$CDE]=\frac{CE\cdot CD}{CA\cdot CB}[ABC].$
Figure 21.
Proof.
This follows from the area formulas $CDE]=\frac{1}{2}CD\cdot CE\sin C$ and $ABC]=\frac{1}{2}CA\cdot CB\sin C$
Theorem 11.2 (Area of Cevian Triangle).
Let $P$ be a point inside $\triangle ABC$. Let APAP meet BCBC at $D$, BPBP meet CACA at $E$ and CPCP meet ABAB at $F$. Suppose $[ABC]=K$, $AB=c$, $BC=a$, $CA=b$, $BD=a_{1}$, $DC=a_{2}$, $CE=b_{1}$, $EA=b_{2}$, $AF=c_{1}$, and $FB=c_{2}$ as shown in Figure 22. Then
$DEF]=\left(1-\frac{c_{1}b_{2}}{bc}-\frac{a_{1}c_{2}}{ca}-\frac{b_{1}a_{2}}{ac}\right)K.$
Triangle DEFDEF is known as the cevian triangle of point $P$.
Figure 22.
Proof.
From the Common Angle Theorem, we have
$[AFE]$ $=\frac{c_{1}b_{2}}{bc}K,$
$[BDF]$ $=\frac{a_{1}c_{2}}{ca}K,$
$[CED]$ $=\frac{b_{1}a_{2}}{ac}K.$
Since [DEF]=[ABC][AFE]BDF][CED][DEF]=[ABC]-[AFE]-BDF]-[CED], we have
$DEF]=\left(1-\frac{c_{1}b_{2}}{bc}-\frac{a_{1}c_{2}}{ca}-\frac{b_{1}a_{2}}{ac}\right)K$
as required. ∎
Yff [8] gave a formula for the area of the cevian triangle of $Y_{1}$, but he gave no proof. We now give a purely geometric proof of his result.
Theorem 11.3 (Yff Area Formula).
In $\triangle ABC$, let $AY_{1}$ meet BCBC at $D$, $BY_{1}$ meet CACA at $E$ and $CY_{1}$ meet ABAB at $F$ (Figure 23). Then
$DEF]=\frac{u^{3}}{2R}.$
Figure 23.
Proof.
From the definition of $Y_{1}$, the segments along the sides of $\triangle ABC$ have the lengths shown in Figure 23. From Theorem 11.2, we have
$[DEF]$ $=\left(1-\frac{u(b-u)}{bc}-\frac{u(c-u)}{ca}-\frac{u(a-u)}{ab}\right)K$
$=\frac{u^{3}+(a-u)(b-u)(c-u)}{abc}K$
$=\frac{u^{3}+u^{3}}{abc}K$
$=2u^{3}\times\frac{K}{abc}$
$=\frac{u^{3}}{2R}$
since $R=abc/(4K)$. ∎
In a similar manner, we find that the area of the cevian triangle of $Y_{2}$ is also $u^{3}/(2R)$.
Weisstein [7] calls these triangles Yff Triangles.
12. Special Points
Let $Y_{m}$ be the midpoint of $Y_{1}Y_{2}$. The point $Y_{m}$ is a triangle center, but it is not one of the ones cataloged in ETC [5] as of January 1, 2026. The first normalized trilinear coordinate of $Y_{m}$ in a 6–9–13 triangle is 0.4500479513210176.
Let $Y_{c}$ be the center of the unique conic that passes through the points AA, $B$, $C$, $Y_{1}$, and $Y_{2}$. The point $Y_{c}$ is not a triangle center listed in ETC [5] as of January 1, 2026. The first normalized trilinear coordinate of $Y_{c}$ in a 6–9–13 triangle is 0.7440584232553069.